# A mathematical theory of topological invariants of quantum spin systems

### v1

@@ 6.3 The Hall conductance
$\sigma(\beta)=\psi\left(\sum_{i<j}\beta_{ij}[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]\right).$ (110)
@@ 6.4 Topological invariants of gapped states on subsets of ${\mathbb{R}}^{n}$
So far we assumed that $\Lambda$ is an arbitrary uniformly discrete subset of ${\mathbb{R}}^{n}$. Suppose $\Lambda\subset W^{\epsilon}$ for some CS set $W$ and some $\epsilon>0$. If this is the case, we will say the pair $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ describes a quantum lattice system on $W$. Then for any ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ the component ${\mathsf{F}}^{X}$ vanishes for any brick $X$ which does not intersect $W^{\epsilon}\cap\Lambda$, and thus for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ we have ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(U\wedge W)$ and ${\mathfrak{D}}^{\psi}_{al}(U)={\mathfrak{D}}^{\psi}_{al}(U\wedge W)$. Thus the assignment $U\mapsto{\mathfrak{D}}^{\psi}_{al}(U)$ can be regarded as a local Lie algebra over the site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$.
@@
    $(\partial{\mathsf{f}})^{Y}:=\sum_{j\in{\mathbb{Z}}^{n}\cap U}{\mathsf{f}}_{j}^{Y}$ (112)
@@
$1+d(Y,U^{1}\cap{\mathbb{Z}}^{n})+\operatorname{diam}(Y)\leq 2(1+d(Y,U)+\operatorname{diam}(Y))$ (113)
@@
$d(Y,U^{1}\cap{\mathbb{Z}}^{n})$ $\leq\operatorname{diam}(Y)+d(y,z)$ (114)
$\leq\operatorname{diam}(Y)+d(y,u)+d(u,z)$ (115)
$\leq\operatorname{diam}(Y)+d(Y,U)+1,$ (116)
@@
$\|{\mathsf{f}}_{i}\|_{k}$ $=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,\{i\})^{k}\|{\mathsf{f}}_{i}^{Y}\|$ (118)
$=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,U^{1}\cap{\mathbb{Z}}^{1}))^{k}\|{\mathsf{F}}^{Y}\|,$ (119)
@@
$\|\partial{\mathsf{f}}^{Y}\|$ $\leq\sum_{j\in{\mathbb{Z}}^{n}\cap U}\|{\mathsf{f}}_{j}^{Y}\|$ (120)
$\leq\|{\mathsf{f}}\|_{k+2n+1}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-k-2n-1}$ (121)
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-2n-1}$ (122)
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$ (123)
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$ (124)
@@
$\|{\mathsf{a}}\|^{KS}_{x,k}:=\sup_{r>0}(1+r)^{k}\inf_{{\mathsf{b}}\in{\mathfrak{d}}(B_{r}(x))}\|{\mathsf{a}}-{\mathsf{b}}\|$ (125)
@@
$\|{\mathsf{b}}-{\mathsf{a}}\|\leq\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-k-2n-2}$ (128)
@@
$\leq(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-2n-2}$ (129)
$\leq C(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}$ (130)
@@
$\|{\mathsf{a}}^{X}\|$ $=\|({\mathsf{a}}-{\mathsf{b}})^{X}\|$ (131)
$\leq 4^{n}\|{\mathsf{a}}-{\mathsf{b}}\|$ (132)
$\leq 4^{n}\|{\mathsf{a}}\|_{x,k}^{KS}(1+r)^{-k}$ (133)
$\leq 4^{n+2k}\|{\mathsf{a}}\|_{x,k}^{KS}(1+\operatorname{diam}(X)+d(x,X))^{-k},$ (134)
@@ Appendix B Inhomogeneous Maurer-Cartan equation
$\xi_{\mathsf{a}}({\mathsf{p}})=[{\mathsf{a}},{\mathsf{p}}]-\partial{\mathsf{a}},$ (137)
@@
$\partial\tilde{\mathsf{p}}_{N}+\frac{1}{2}[\tilde{\mathsf{p}}_{N},\tilde{\mathsf{p}}_{N}]={\mathsf{B}}_{N}+{\mathsf{q}}_{N}$ (139)
@@
$\exp({\mathsf{b}}+{\mathsf{a}})({\mathsf{p}})=\exp({\mathsf{b}})({\mathsf{p}})-\partial{\mathsf{a}}.$ (140)
@@
$\partial{\mathsf{q}}=\frac{1}{2}[{\mathsf{p}}_{2},{\mathsf{p}}_{2}]-\frac{1}{2}[{\mathsf{p}}_{1},{\mathsf{p}}_{1}]=[{\mathsf{q}},{\mathsf{p}}_{1}]+\frac{1}{2}[{\mathsf{q}},{\mathsf{q}}]=0.$ (141)
@@
${\mathsf{p}}^{\prime}-{\mathsf{p}}=-\partial{\mathsf{a}}+\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!},$ (142)
@@
${\mathsf{f}}:=\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!}\in[{\mathfrak{M}},{\mathfrak{M}}].$ (143)
Thus $[{\mathsf{p}},{\mathsf{p}}]-[{\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]=2\partial({\mathsf{p}}^{\prime}-{\mathsf{p}})=2\partial{\mathsf{f}}$ is $\partial$-exact in $[{\mathfrak{M}},{\mathfrak{M}}]$. ∎

### v2

@@ 6.3 The Hall conductance
$\sigma(\beta)=\psi\left(\sum_{i<j}\beta_{ij}[{\mathsf{Q}}_{i},{\mathsf{Q}}_{j}]\right).$
@@ 6.4 Topological invariants of gapped states on subsets of ${\mathbb{R}}^{n}$
So far we assumed that $\Lambda$ is an arbitrary countable subset of ${\mathbb{R}}^{n}$ with a uniform $O(r^{n})$ bound on the number of points in a ball of radius $r$. Suppose $\Lambda\subset W^{\epsilon}$ for some CS set $W$ and some $\epsilon>0$. If this is the case, we will say the pair $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ describes a quantum lattice system on $W$. Then for any ${\mathsf{F}}\in{\mathfrak{D}}_{al}$ the component ${\mathsf{F}}^{X}$ vanishes for any brick $X$ which does not intersect $W^{\epsilon}\cap\Lambda$, and thus for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ we have ${\mathfrak{D}}_{al}(U)={\mathfrak{D}}_{al}(U\wedge W)$ and ${\mathfrak{D}}^{\psi}_{al}(U)={\mathfrak{D}}^{\psi}_{al}(U\wedge W)$. Thus the assignment $U\mapsto{\mathfrak{D}}^{\psi}_{al}(U)$ can be regarded as a local Lie algebra over the site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$.
@@
    $(\partial{\mathsf{f}})^{Y}:=\sum_{j\in{\mathbb{Z}}^{n}\cap U}{\mathsf{f}}_{j}^{Y}$
@@
$1+d(Y,U^{1}\cap{\mathbb{Z}}^{n})+\operatorname{diam}(Y)\leq 2(1+d(Y,U)+\operatorname{diam}(Y))$
@@
$d(Y,U^{1}\cap{\mathbb{Z}}^{n})$ $\leq\operatorname{diam}(Y)+d(y,z)$
$\leq\operatorname{diam}(Y)+d(y,u)+d(u,z)$
$\leq\operatorname{diam}(Y)+d(Y,U)+1,$
@@
$\|{\mathsf{f}}_{i}\|_{k}$ $=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,\{i\})^{k}\|{\mathsf{f}}_{i}^{Y}\|$
$=\sup_{Y\in{\mathbb{B}}_{n}\backslash\{\varnothing\}}(1+\operatorname{diam}(Y)+d(Y,U^{1}\cap{\mathbb{Z}}^{1}))^{k}\|{\mathsf{F}}^{Y}\|,$
@@
$\|\partial{\mathsf{f}}^{Y}\|$ $\leq\sum_{j\in{\mathbb{Z}}^{n}\cap U}\|{\mathsf{f}}_{j}^{Y}\|$
$\leq\|{\mathsf{f}}\|_{k+2n+1}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-k-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}\sum_{j\in{\mathbb{Z}}^{n}\cap U}(1+\operatorname{diam}(Y)+d(j,Y))^{-2n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$
$\leq\|{\mathsf{f}}\|_{k+2n+1}(1+\operatorname{diam}(Y)+d(U,Y))^{-k}(1+\operatorname{diam}(Y))^{-n}\sum_{j\in{\mathbb{Z}}^{n}}(1+d(j,Y))^{-n-1}$
@@
$\|{\mathsf{a}}\|^{KS}_{x,k}:=\sup_{r>0}(1+r)^{k}\inf_{{\mathsf{b}}\in{\mathfrak{d}}(B_{r}(x))}\|{\mathsf{a}}-{\mathsf{b}}\|$
@@
$\|{\mathsf{b}}-{\mathsf{a}}\|\leq\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-k-2n-2}$
@@
$\leq(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}\sum_{X\nsubseteq B_{r}(x)}(1+\operatorname{diam}(X)+d(x,X))^{-2n-2}$
$\leq C(1+r)^{-k}\|{\mathsf{a}}\|_{\{x\},k+2n+2}$
@@
$\|{\mathsf{a}}^{X}\|$ $=\|({\mathsf{a}}-{\mathsf{b}})^{X}\|$
$\leq 4^{n}\|{\mathsf{a}}-{\mathsf{b}}\|$
$\leq 4^{n}\|{\mathsf{a}}\|_{x,k}^{KS}(1+r)^{-k}$
$\leq 4^{n+2k}\|{\mathsf{a}}\|_{x,k}^{KS}(1+\operatorname{diam}(X)+d(x,X))^{-k},$
@@ Appendix B Inhomogeneous Maurer-Cartan equation
$\xi_{\mathsf{a}}({\mathsf{p}})=[{\mathsf{a}},{\mathsf{p}}]-\partial{\mathsf{a}},$
@@
$\partial\tilde{\mathsf{p}}_{N}+\frac{1}{2}[\tilde{\mathsf{p}}_{N},\tilde{\mathsf{p}}_{N}]={\mathsf{B}}_{N}+{\mathsf{q}}_{N}$
@@
$\exp({\mathsf{b}}+{\mathsf{a}})({\mathsf{p}})=\exp({\mathsf{b}})({\mathsf{p}})-\partial{\mathsf{a}}.$
@@
$\partial{\mathsf{q}}=\frac{1}{2}[{\mathsf{p}}_{2},{\mathsf{p}}_{2}]-\frac{1}{2}[{\mathsf{p}}_{1},{\mathsf{p}}_{1}]=[{\mathsf{q}},{\mathsf{p}}_{1}]+\frac{1}{2}[{\mathsf{q}},{\mathsf{q}}]=0.$
@@
${\mathsf{p}}^{\prime}-{\mathsf{p}}=-\partial{\mathsf{a}}+\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!},$
@@
${\mathsf{f}}:=\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}({\mathsf{p}})}{k!}-\sum_{k=1}^{\infty}\frac{{\rm ad}^{k}_{\mathsf{a}}(\partial{\mathsf{a}})}{(k+1)!}\in{\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}.$
Thus $[{\mathsf{p}},{\mathsf{p}}]-[{\mathsf{p}}^{\prime},{\mathsf{p}}^{\prime}]=2\partial({\mathsf{p}}^{\prime}-{\mathsf{p}})=2\partial{\mathsf{f}}$ is $\partial$-exact in ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$. ∎